# Fragmented eigenstate thermalization versus robust integrability in long-range models

### v1

@@ I Justification on why tuning $\alpha$ is a non-local perturbation to $\hat{H}^{\alpha=0}$
$\epsilon(\alpha)=\frac{||\hat{V}_{\alpha}||}{||\hat{H}^{\alpha=0}||},$ (S7)
where the Hilbert-Schmidt norm is given by $||A||=\sqrt{\sum_{i}\lambda_{i}^{2}}$, with $\lambda_{i}$ denoting the eigenvalues of the Hermitian matrix $A$. As shown in Fig. S4, we find $\epsilon(\alpha)<1$ for $\alpha<1$, validating the interpretation of $\hat{V}_{\alpha}$ as a perturbation to $\hat{H}^{\alpha=0}$.
Figure S4: For the Hamiltonian given in Eq. (1) of the main text with $\alpha=0$, we demonstrate that introducing a small but finite $\alpha$ acts as a perturbation, as the ratio of the norms can be fitted to the form $\epsilon=0.49\,\alpha^{1.0}$.
II Discussion on the total spin conservation at $\alpha=0$: Band structures in DOS
We begin by recalling Eq. (2) of the main text, which corresponds to the $\alpha=0$ limit of the Hamiltonian in Eq. (1):
$\hat{H}^{\alpha=0}=\frac{2J}{N}\hat{S}_{x}^{2}+2h\hat{S}_{z}-2J,$ (S8)
@@ II Discussion on the total spin conservation at $\alpha=0$: Band structures in DOS
[…]
@@
[…]
Figure S5: For the fully connected limit of the LRTFIM with the Hamiltonian given in Eq. (S8) for $L=8$ spins, panel (a) shows the density of states (DOS), revealing 25 distinct energy bands. Panel (b) confirms that $\hat{S}^{2}$ is diagonal in the computational eigenbasis of Eq. (S8), and panel (c) illustrates how each band can be uniquely identified by its corresponding total spin quantum number $s$.
Another important aspect is to determine for a given $L$, if one can find the maximum $s$-sector. This boils down to the task of maximizing the function $n(s,L)$ with respect to $s$ for a given $L$. This can be easily numerically settled by just observing that the maximal sector (the sector that has most number of eigenstates in each band) $s_{\mathrm{max}}\equiv\max_{s}{n(s,L)}$ varies as a function of $L$, as demonstrated in Fig. S6. Identifying the $s_{\mathrm{max}}$ is particularly relevant as one is usually interested in understanding ETH or thermalization dynamics in the most populated symmetry sector.
Figure S6: For the Hamiltonian at $\alpha=0$ in Eq. (S8), we present the characterization of maximally populated $s_{\mathrm{max}}$-sector for increasing $L$. Even for numerically accessible system sizes, we find that the variation of $s_{\mathrm{max}}$ with $L$ is quite relevant.
@@ III Degenerate perturbation theory: Why a certain class gives chaos while others retain integrability?
In this section, we provide a first-principles explanation for the emergence of either integrability or chaos in the cases of class (i), class (ii), and class (iii) perturbations. To this end, we employ degenerate perturbation theory with the setup
@@
where $\hat{H}^{\alpha=0}$ is the unperturbed Hamiltonian as given in Eq. (2) of the main text, and $\hat{V}$ denotes the perturbation belonging to one of the classes discussed previously. As outlined earlier, the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$, owing to its $SU(2)$ symmetry, exhibits a highly degenerate density of states (DOS), organized into bands labeled by the total spin quantum number $s$. The method for analyzing the effect of perturbations is as follows:
  • First, we pin down the symmetries of the Hamiltonian $\hat{H}$ (perturbation included) and we numerically obtain the spectrum of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$ in a symmetry-resolved manner.
  • Next, we numerically identify the most populated energy band in the DOS of $\hat{H}^{\alpha=0}$, with energy $E_{0}$, and restrict our analysis to the corresponding eigenstates $\{|m_{0}\rangle\}$ within this band.
  • For each pair of states $|m_{0}\rangle$, $|n_{0}\rangle$ in this band, we compute the matrix elements of the perturbation: $\langle m_{0}|\hat{V}|n_{0}\rangle$.
@@
If most of the eigenvalues $\lambda_{i}$ vanish, the degeneracy is largely preserved, indicating integrability. If the eigenvalues are non-zero but exhibit degeneracies or clustering, the original band splits into multiple subbands, and the system retains integrable features. On the other hand, if the degeneracy is completely lifted, one must analyze the statistics of the corrected energy levels: the presence of level repulsion indicates the emergence of chaos, while level attraction or clustering suggests persistent integrability.
To proceed, we consider the three classes separately.
@@ III.A Class (i): Non-extensive
One-body perturbations. We introduce a single-site impurity at a bulk site, either in the longitudinal or transverse direction. For a longitudinal impurity, the perturbation takes the form $\delta\hat{\sigma}^{x}_{L/2}$. In this case, we find that all first-order corrections vanish, i.e., $\lambda_{n_{0}}^{i}=0$ for all $i$, indicating that the degeneracy of the unperturbed band remains completely intact. In contrast, a transverse impurity, given by $\delta\hat{\sigma}^{z}_{L/2}$, lifts the degeneracy: all eigenvalues $\lambda_{i}$ become nonzero. However, the spectrum of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibits internal degeneracies, leading to a splitting of the original degenerate band into several subbands, each with its own residual degeneracy.
Although perturbation theory is strictly valid only in the $\delta\ll 1$ regime, we observe that this qualitative structure persists even for larger values of $\delta$. These results explain the persistence of integrability in the presence of such local perturbations. We also note that in the other integrable limit $\alpha\to\infty$, a longitudinal impurity of the same kind would instead induce chaos, in stark contrast to the behavior observed here.
Two-body perturbations. In the bulk, we consider a single two-body interaction term, either in the longitudinal or transverse direction, given respectively by $\hat{\sigma}^{x}_{L/2-1}\hat{\sigma}^{x}_{L/2}$ or $\hat{\sigma}^{z}_{L/2-1}\hat{\sigma}^{z}_{L/2}$. Similar to the one-body case, the eigenvalues $\{\lambda_{n_{0}}^{i}\}$ of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibit degeneracies, indicating that a single band splits into multiple subbands, each retaining some degeneracy. This spectral structure reflects the persistence of integrability despite the perturbation. We also note that the choice of the specific pair $(L/2-1,L/2)$ is not essential; any arbitrary pair in the bulk produces qualitatively similar results.
@@ III.B Class (ii): Extensive one-body
For homogeneous perturbations as in the form of $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$ or $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}$, the perturbed Hamiltonian still retains the original $SU(2)$ symmetry by conserving the total spin $\hat{S}^{2}$. This leads to the organization of the spectrum into more degenerate bands, thereby trivially extending the integrability of the unperturbed system.
For inhomogeneous perturbations in the transverse direction, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{z}$, with $h_{i}\in[-\delta,\delta]$ being random, we find that the degeneracies are getting lifted inside a band. Also, the level-ratio test reveals a clear Poissonian signature, implying integrability. Although the original $SU(2)$ conservation is lost, the first-order corrections reveal zero correlation, which we understand as a hint towards integrability for exact energy statistics. For random field perturbations in the longitudinal direction $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}$, the situation is trickier, and a discussion on its exact eigenvalue statistics, as well as the behaviour of the spectral form factor, and matrix elements inside a band for infinitesimal perturbation is presented in the main text.
@@ III.C Class (iii): Extensive two-body
In this class of perturbations, we have nearest-neighbour (nn) and interactions beyond nn, evidently incorporating the perturbation in $\alpha$ by turning on infinitesimal $\alpha$ from 0.
Let us first consider the case when the perturbation is nearest-neighbor and given by
@@
We begin by examining the level statistics in the central one-third of the spectrum, resolved by both parity and inversion symmetries. Even for very small values of $\delta$, we observe Wigner–Dyson statistics, indicative of quantum chaos.
@@
To further investigate the mechanism underlying this chaotic behavior, we construct the matrix $\langle m^{0}|\hat{V}^{\text{nn}}|n^{0}\rangle$ within the most populated degenerate band of $\hat{H}^{\alpha=0}$, restricted to the relevant symmetry sector. Diagonalizing this matrix, we obtain the first-order energy corrections $\{\lambda_{n_{0}}^{i}\}$. All the eigenvalues are found to be distinct, indicating that the degeneracy is completely lifted at first order in perturbation theory.
Remarkably, even the set of first-order corrections $\{\lambda_{n_{0}}^{i}\}$ exhibits level repulsion: the distribution of the unfolded consecutive level spacings $\delta\lambda_{n}^{i}=\lambda_{{n+1}_{0}}^{i}-\lambda_{n_{0}}^{i}$, after sorting the eigenvalues energetically, follows Wigner’s surmise. The emergence of level repulsion (see Fig. S7) already at the level of the first-order corrections provides a clear mechanism for the onset of chaos in the perturbed system.

### v2

@@ I Why tuning $\alpha$ is a non-local perturbation to $\hat{H}^{\alpha=0}$
where the Hilbert-Schmidt norm is given by $A||=\sqrt{\sum_{i}\lambda_{i}^{2}}$, with $\lambda_{i}$ denoting the eigenvalues of the Hermitian matrix AA. As shown in Fig. S3, we find $\varepsilon(\alpha)<1$ for $\alpha<1$, validating the interpretation of $\hat{V}_{\alpha}$ as a perturbation to $\hat{H}^{\alpha=0}$.
Figure S3: For the Hamiltonian given in Eq. (1) of the main text with $\alpha=0$, this figure demonstrates that introducing a small but finite $\alpha$ acts as a perturbation, since the ratio of the norms can be fitted to the form $\varepsilon=0.49\,\alpha^{1.0}$.
II Total spin conservation at $\alpha=0$: Density of states Band structures
The Hamiltonian in Eq. (2) of the main text, corresponding to the $\alpha=0$ limit of the Hamiltonian in Eq. (1) of the main text, presents $SU(2)$ symmetry, which implies conservation of the total spin, $\hat{S}^{2},\hat{H}^{\alpha=0}]=0$, with $\hat{S}^{2}\equiv\hat{S}_{x}^{2}+\hat{S}_{y}^{2}+\hat{S}_{z}^{2}$. This symmetry leads to a highly degenerate spectrum, with the density of states exhibiting distinct energy bands, each containing a large number of degenerate eigenstates.
Since $\hat{S}^{2}$ is a symmetry of the Hamiltonian, one can characterize each energy band (namely the eigenstates in the band) by its total spin value $s$. For a given $L$ (taken to be even), the total spin can take values $s=0,1,\ldots,L/2$. Due to the presence of the transverse field, each $s$-sector further splits into $(2s+1)$ bands. Therefore, for a given $L$, the total number of bands, each corresponding to a distinct energy eigenvalue, is given by
@@ II Total spin conservation at $\alpha=0$: Density of states Band structures
The number of eigenstates in each energy band depends only on the corresponding total spin quantum number $s$. This number corresponds to the degeneracy associated with total spin resulting from the addition of angular momenta of $L$ spin-$\tfrac{1}{2}$ particles. Following the standard rules encoded in Catalan’s triangle for combining spin-$\tfrac{1}{2}$ particles, we obtain that for any $s<L/2$, the number of eigenstates in each of the $(2s+1)$ bands is given by
@@
whereas for $s=L/2$, corresponding to the fully symmetric sector, there is no degeneracy, so each of the $(L+1)$ eigenstates is non-degenerate.
In Fig. S4, we illustrate the band structure of the density of states (DOS) for $L=8$, along with the corresponding characterization of each band by the total spin quantum number $s$.
Figure S4: For the fully connected limit of the long-range transverse field Ising model with the Hamiltonian given in Eq. (2) of the main text and $L=8$ spins, panel (a) shows the density of states (DOS), revealing 25 distinct energy bands. Panel (b) confirms that $\hat{S}^{2}$ is diagonal in the eigenstates of the Hamiltonian in Eq. (2) of the main text. Panel (c) shows that each band is uniquely identified by its corresponding total spin quantum number $s$.
An additional key point is to identify, for a given system size $L$, the sector $s$ that contains the largest number of states. Determining this $s_{\mathrm{max}}$ is important, because studies of quantum chaos and ETH typically focus on the most populated symmetry sector. This tasks amounts to maximizing $n(s,L)$ with respect to $s$, which can be straightforwardly addressed numerically. As shown in Fig. S5, the size of the most populated sector, $s_{\mathrm{max}}\equiv\max_{s}n(s,L)$, varies as a function of $L$.
Figure S5: For the Hamiltonian at $\alpha=0$ in Eq. (2) of the main text, we present the maximally populated $s_{\mathrm{max}}$-sector for increasing $L$. Even for numerically accessible system sizes, we find variations of $s_{\mathrm{max}}$ with $L$.
@@ III Degenerate perturbation theory: Why a certain class gives chaos while others retain integrability?
In this section, we provide a first-principles explanation for the emergence of either integrability or chaos in the cases of (i) non-extensive, (ii) extensive one-body, and (iii) extensive two-body perturbations. To this end, we employ degenerate perturbation theory with the setup
@@
where $\hat{H}^{\alpha=0}$ is the unperturbed Hamiltonian as given in Eq. (2) of the main text, and $\hat{V}$ denotes the perturbation belonging to one of the three classes above. The method for analyzing the effect of the perturbations is as follows:
  • First, we pin down the symmetries of the Hamiltonian $\hat{H}$ (perturbation included) and numerically obtain the spectrum of the unperturbed Hamiltonian $\hat{H}^{\alpha=0}$ in a symmetry-resolved manner.
  • Next, we numerically identify the most populated energy band in the DOS of $\hat{H}^{\alpha=0}$, with energy $E_{n_{0}}$, and restrict our analysis to the corresponding eigenstates $\{|n_{0}\rangle\}$ within this band.
  • For each pair of states $|m_{0}\rangle$, $|n_{0}\rangle$ in this band, we compute the matrix elements of the perturbation, $\langle m_{0}|\hat{V}|n_{0}\rangle$.
@@
If most of the eigenvalues $\lambda^{i}_{n_{0}}$ vanish, the degeneracy is largely preserved, indicating integrability. If the eigenvalues are non-zero but exhibit degeneracies or clustering, the original band splits into multiple subbands, and the system may retain integrable features. On the other hand, if the degeneracy is lifted, the analysis of level statistics determines whether the system is chaotic (level repulsion) or not. In the next subsections, we consider the three classes separately.
@@ III.A Class (i): Non-extensive
One-body perturbations. We introduce a single-site impurity at a bulk site, either in the longitudinal or transverse direction. For a longitudinal impurity, the perturbation takes the form $\delta\hat{\sigma}^{x}_{L/2}$. In this case, we find that all first-order corrections vanish, i.e., $\lambda_{n_{0}}^{i}=0$ for all $i$, indicating that the degeneracy of the unperturbed band remains intact. In contrast, a transverse impurity, given by $\delta\hat{\sigma}^{z}_{L/2}$, leads to nonzero values of $\lambda^{i}_{n_{0}}$. However, the spectrum of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibits internal degeneracies, leading to a splitting of the original degenerate band into several subbands, each with its own residual degeneracy.
Although perturbation theory is strictly valid only for $\delta\ll 1$, we observe that this qualitative structure persists even for larger values of $\delta$. These results explain the persistence of integrability in the presence of such local perturbations. We also note that in the other integrable limit $\alpha\to\infty$, a longitudinal impurity of the same kind would instead induce chaos, in stark contrast to the behavior observed here.
Two-body perturbations. In the bulk, we consider a single two-body interaction term, either in the longitudinal or transverse direction, given respectively by $\hat{\sigma}^{x}_{L/2-1}\hat{\sigma}^{x}_{L/2}$ or $\hat{\sigma}^{z}_{L/2-1}\hat{\sigma}^{z}_{L/2}$. Similar to the one-body case, the eigenvalues $\{\lambda_{n_{0}}^{i}\}$ of the perturbation matrix $\langle m_{0}|\hat{V}|n_{0}\rangle$ exhibit degeneracies, indicating that a single band splits into multiple subbands, each retaining some degree of degeneracy. This spectral structure reflects the persistence of integrability despite the perturbation. We also note that the choice of the specific pair (L/21,L/2)(L/2-1,L/2) is not essential; any arbitrary pair in the bulk produces qualitatively similar results.
Figure S6: Analysis of (a)-(b) spectral correlations and (c) off-diagonal ETH for the most populated energy band of the perturbed Hamiltonian in Eq. (2) of the main text with $\hat{V}=h_{i}\sigma_{i}^{x}$ , random numbers $h_{i}\in[-\delta,\delta]$ and $\delta=10^{-4}$. (a) Distribution of the mean value of the ratio of consecutive level spacings, $\bar{r}$, for various disorder realizations and their average vs system size $L$ in the inset. (b) Spectral form factor for different system sizes; dashed line indicates the saturation value. (c) Distribution of the off-diagonal elements of $\hat{S}_{z}$ for 200 levels in the middle of the energy band, $L=13$. The solid line represents a log-normal distribution.
@@ III.B Class (ii): Extensive one-body
For homogeneous perturbations, as in the form of $\sum_{i=1}^{L}\hat{\sigma}_{i}^{x}$ or $\sum_{i=1}^{L}\hat{\sigma}_{i}^{z}$, the perturbed Hamiltonian still retains the original $SU(2)$ symmetry by conserving the total spin $\hat{S}^{2}$. This leads to the organization of the spectrum into more degenerate bands, thereby trivially extending the integrability of the unperturbed system.
For inhomogeneous perturbations in the transverse direction, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{z}$, with $h_{i}\in[-\delta,\delta]$ being random, we find that the degeneracies are lifted inside a band. Nevertheless, the level statistics analysis reveals a clear Poissonian signature, implying integrability. Although the original $SU(2)$ conservation is lost, the first-order corrections do not reveal spectral correlations, which suggests integrability for the exact energy levels as well.
For random field perturbations in the longitudinal direction, $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{x}$, the situation is more subtle. The situation becomes more intricate with inhomogeneous fields. Transverse inhomogeneous fields, for instance $\sum_{i}h_{i}\hat{\sigma}^{z}_{i}$, break symmetry and lift degeneracies, yet for either weak ($\delta\ll J$) or strong ($\delta\sim{\cal O}(J)$) disorder, we observe Poissonian level statistics for $\lambda_{n_{0}}$, indicating preserved integrability. We speculate that such behaviour is possibly related to Richardson-Gaudin integrability, as we demonstrate later that to be the case for fully-connected XXZXXZ Hamiltonian, where a $\sum_{i=1}^{L}h_{i}\hat{\sigma}_{i}^{z}$ perturbation allows one to construct an extensive set of conserved charges.
However, longitudinal inhomogeneous fields, $\sum_{i}h_{i}\hat{\sigma}^{x}_{i}$, lead to level statistics for $\lambda_{n_{0}}$ intermediate regime between Poisson and Wigner-Dyson distributions, suggesting level repulsion in the presence of possible (quasi-)symmetries. To further assess level repulsion and ETH in this setting, we return to the exact eigenvalues of the Hamiltonian in Eq. (2) of the main text with $\sum_{i}h_{i}\hat{\sigma}^{x}_{i}$. We consider very weak random perturbations ($\delta=10^{-4}$) and analyze in Fig. S6 different indicators of quantum chaos for the most populated energy band. The mean ratio of consecutive level spacings, $\bar{r}$, exhibits strong fluctuations across individual disorder realizations, as indicated by the relatively broad distributions in Fig. S6(a). However, the ensemble average, $\langle\bar{r}\rangle$, increases with system size $L$, as evidenced in Fig. S6(a) by a systematic shift toward larger values and narrowing of the distributions as $L$ increases. The inset confirms this trend, showing that $\langle\bar{r}\rangle$ grows with $L$.
In Fig. S6(b) we examine the spectral form factor, defined as [50]
$K(t)=\left|\frac{1}{\mathcal{D}}\sum_{n=1}^{\mathcal{D}}e^{iE_{n}t}\right|^{2},$ (S11)
where $\mathcal{D}$ is the Hilbert space dimension of the most populated energy band. This quantity sensitively captures spectral correlations by developing a dip–ramp–plateau structure (correlation hole) even in the presence of symmetries [62, 21]. The correlation hole, corresponding to the time interval with values of $K(t)$ below its saturation (dashed) line, is seen in Fig. S6(b) for all three system sizes considered, with no sign of reduction of the relative depth with $L$.
Figures S6(a)-(b) suggest that integrability is broken for arbitrarily weak random longitudinal fields, yet quantum chaos is not fully developed. This is further supported by Fig. S6(c), where off-diagonal ETH is probed with the distribution of the matrix elements of $\hat{S}_{z}$, as in Fig. 2(g). Instead of the Gaussian form expected in many-body quantum chaos, we observe a log-normal distribution, indicative of structured eigenstates potentially shaped by residual (quasi-)symmetries. These results point to the random longitudinal field perturbation as a subtle and nontrivial case deserving further theoretical and numerical exploration.
@@ III.C Class (iii): Extensive two-body
In this class of perturbations, we can consider nearest-neighbor interactions and beyond, as when incorporating the perturbation in $\alpha$ by changing it infinitesimally from 0. The case $\alpha\rightarrow 0$ is shown in the main text. Let us consider here the case when the perturbation is nearest-neighbor and given by
@@
We first examine the level statistics in the central one-third of the spectrum, resolved by both parity and inversion symmetries. Even for very small values of $\delta$, we observe Wigner–Dyson statistics, indicative of quantum chaos. To further investigate the mechanism underlying this chaotic behavior, we construct the matrix $\langle m_{0}|\hat{V}^{\text{nn}}|n_{0}\rangle$ within the most populated degenerate band of $\hat{H}^{\alpha=0}$, restricted to the relevant symmetry sector. Diagonalizing this matrix, we obtain the first-order energy corrections $\{\lambda_{n_{0}}^{i}\}$. All the eigenvalues are found to be distinct, indicating that the degeneracy is completely lifted at first order in perturbation theory. Even the set of first-order corrections $\{\lambda_{n_{0}}^{i}\}$ exhibits level repulsion. The distribution of the unfolded consecutive level spacings $\delta\lambda_{n}^{i}=\lambda_{{n+1}_{0}}^{i}-\lambda_{n_{0}}^{i}$, after sorting the eigenvalues energetically, follows the Wigner-Dyson distribution. The emergence of level repulsion (see Fig. S7) already at the level of the first-order corrections provides a clear mechanism for the onset of chaos in the perturbed system. We note that the same holds for all perturbations in this class.